Donna invested her $33,000 bonus and reieved a tatal of $970 in interest after one year. If part of the money returned 4% and the remainder 2.25% then how much did she invest at each rate?
THANK U :)

Answer by checkley71(8403)

x*.04+(33000-x)*.0225=970 or.04x+742.5-.0225X=970 or .0175x=970-742.5 or
.0175x=227.5 or x=227.5/.0175 or x=13000 invested @4% the 33000-13000=20000
invested @2.25%